OpenStudy (anonymous):

@lulu22 A credit card had an APR of 11.91% all of last year and compounded interest daily. What was the credit card's effective interest rate last year? A. 11.91% B. 13.37% C. 12.64% D. 17.84%

OpenStudy (amistre64):

isnt effective yeild a term for the APY?

OpenStudy (amistre64):

\[(1+\frac{apr}{n})^n=(1+apy)\] \[(1+\frac{apr}{n})^n-1=apy\]

OpenStudy (amistre64):

since there are 365 days in a year, n = 365 to run that calculations with